A high school with 425 students (16% African-American) closes. All of the students join a high school whose student body of 1250 students was 44% African-American. Rounded to the nearest per cent, the per cent of African-American students in the high school after the merger is

A. 30% B 34% C. 37% D. 49 E. 60%

Answer by Theo(3464)

high school with 425 students has .16 8 425 = 68 african americans in it.
high school with 1250 students has .44 * 1250 = 550 african americans in it.

schools merge.

total of students in merged school is 1250 + 425 = 1675.

total of african americans in merged school is 68 + 550 = 618.

percent african americans in combined school is 618 / 1675 = .368955224 = 37%.

your answer is selection C.